# Deep Signature: Characterization of Large-Scale Molecular Dynamics

### v1

@@ Abstract
Understanding protein dynamics are essential for deciphering protein functional mechanisms and developing molecular therapies. However, the complex high-dimensional dynamics and interatomic interactions of biological processes pose significant challenge for existing computational techniques. In this paper, we approach this problem for the first time by introducing Deep Signature, a novel computationally tractable framework that characterizes complex dynamics and interatomic interactions based on their evolving trajectories. Specifically, our approach incorporates soft spectral clustering that locally aggregates cooperative dynamics to reduce the size of the system, as well as signature transform that collects iterated integrals to provide a global characterization of the non-smooth interactive dynamics. Theoretical analysis demonstrates that Deep Signature exhibits several desirable properties, including invariance to translation, near invariance to rotation, equivariance to permutation of atomic coordinates, and invariance under time reparameterization. Furthermore, experimental results on three benchmarks of biological processes verify that our approach can achieve superior performance compared to baseline methods.
@@ 1 Introduction
To this end, we aim to develop a computationally efficient framework that incorporates the structural bioinformatics with coarse graining mapping for automatically analyzing protein trajectory dynamics. In particular, we first introduce a graph clustering module that learns to extract coarse grained dynamics by approximating soft spectral clustering. With the clustering assignment function implemented by a graph neural network and parameters learned automatically, we circumvent the need for manual selection of coarse graining mapping. Subsequently, we introduce a path signature transform module served as a feature extractor to characterize the interatomic interactive dynamics after coarse graining. Path signature is a mathematically principled concept that utilizes iterated integrals to describe geometric rough paths in a compact yet rich manner (Lyons, 2014), thus suitable for our tasks where molecular trajectories are highly sampled and non-smooth. After attaching with a task-specific differentiable classifier or regressor, we devise an end-to-end framework, named Deep Signature, for efficiently characterizing the complex protein dynamics. Notably, due to the existence of considerable random fluctuation in simulated trajectories, ideal features ought to maintain symmetry respecting certain geometric transformations. We provide theoretical analysis that our extracted features exhibit invariance to translation, near invariance to rotation, equivariance to permutation of atomic coordinates, and invariance under time reparametrization of paths. Finally, we target our task on predicting functional properties of proteins from MD data, a fundamental task for developing novel drug therapies. We consider three benchmarks including gene regulatory dynamics, epidermal growth factor receptor (EGFR) mutation dynamics, and G protein-coupled receptors (GPCR) dynamics for performance evaluation. In summary, we make the following three key contributions:
@@ 2 Related Works
The current investigation into characterizing molecular dynamics, especially interatomic interaction dynamics, remains limited, with only a few studies close to ours. Among them, Endo et al. (2019), Yasuda et al. (2022) and Mustali et al. (2023) are unsupervised methods that build local dynamics ensembles for manually selected particles and inspect each particle’s contribution to the dynamics independently. Li et al. (2022) proposes to convert MD conformations into images and employ convolutional neural network to identify diverse active states from MD trajectories. Nevertheless, these approaches cannot recognize subtle interatomic interactions among atomic pathways for dissecting protein function, nor is the complexity of their representations independent of time stamps.
@@ 3.2 Deep Signature
$\mathrm{LogSig}^{D}_{r_{i},r_{i+1}}(\tilde{\mathbf{X}})=\sum_{n=1}^{3M}S(\tilde{\mathbf{X}})^{n}_{r_{i},r_{i+1}}e_{n}+\sum_{1\leq n<m\leq 3M}\frac{1}{2}(S(\tilde{\mathbf{X}})^{n,m}_{r_{i},r_{i+1}}-S(\tilde{\mathbf{X}}))^{m,n}_{r_{i},r_{i+1}}[e_{n},e_{m}]+\dots$ (10)
@@ 4.1 Experiments on Gene regulatory dynamics
For gene regulatory dynamics, we generate 100 trajectories that describe the interactive dynamics between genes and transcription factors. These dynamics are categorized into degradation type or dimerization type, with an equal number of trajectories for each type. The systems are simulated over a period of $2$s and with a time interval of $0.004$s, resulting in 500 frames per trajectory. Besides, each system encompasses 100 nodes with randomly generated Power-law network structure to describe their relationships. We implement the deep spectral clustering module in our approach using two graph pooling layers that first coarsen the dynamics into 60 nodes and then into 30 nodes. The coefficients of loss terms are set as $\lambda_{1}=1$, $\lambda_{2}=0.01$, and $\lambda_{3}=10$. We optimize our model using Adam with an initial learning rate of 5e-4 and a weight decay of 1e-4. For comparison, we consider several baseline approaches which aggregate nodes’ dynamics by averaging without taking into account their interactions. These methods incorporate a deep spectral clustering model with an MLP to process the first frame only (Head), the last frame only (Tail), the first and last frame (Head & Tail), and all frames with an additional LSTM layer to aggregate nodes’ temporal dynamics (GraphLSTM). The overall comparison results with baseline methods are presented in Table 4.
@@
Head 55.280 0.306
Tail 55.360 0.307
Head & Tail 67.920 0.358
GraphLSTM 96.400 0.931
Deep Signature 99.120 0.986
@@ 4.2 Experiments on epidermal growth factor receptor mutation dynamics
$\times$ $\times$ $\checkmark$ 67.600 0.196
$\times$ $\checkmark$ $\checkmark$ 67.967 0.176
$\checkmark$ $\times$ $\checkmark$ 66.000 0.232
$\checkmark$ $\checkmark$ $\checkmark$ 69.333 0.220
@@ 4.3 Experiments on G protein-coupled receptors dynamics
Accuracy 44.000 50.000 46.667 45.467 48.533 53.333 64.200
Recall 0.355 0.667 0.366 0.366 0.385 0.360 0.413
[…]
@@ B.1 Gene regulatory dynamics
$\frac{\mathrm{d}\bm{x}_{t}(v_{i})}{\mathrm{d}t}=-b_{i}\bm{x}(v_{i})f+\sum_{j=1}^{n}\mathbf{A}^{(i,j)}\frac{\bm{x}^{h}(v_{j})}{\bm{x}^{h}(v_{j})+1},$ (13)

### v2

@@ Abstract
Understanding protein dynamics are essential for deciphering protein functional mechanisms and developing molecular therapies. However, the complex high-dimensional dynamics and interatomic interactions of biological processes pose significant challenge for existing computational techniques. In this paper, we approach this problem for the first time by introducing Deep Signature, a novel computationally tractable framework that characterizes complex dynamics and interatomic interactions based on their evolving trajectories. Specifically, our approach incorporates soft spectral clustering that locally aggregates cooperative dynamics to reduce the size of the system, as well as signature transform that collects iterated integrals to provide a global characterization of the non-smooth interactive dynamics. Theoretical analysis demonstrates that Deep Signature exhibits several desirable properties, including invariance to translation, near invariance to rotation, equivariance to permutation of atomic coordinates, and invariance under time reparameterization. Furthermore, experimental results on three benchmarks of biological processes verify that our approach can achieve superior performance compared to baseline methods. Our code is available at https://github.com/WonderSeven/Deep-Signature.
@@ 1 Introduction
To this end, we aim to develop a computationally efficient framework that incorporates the structural bioinformatics with coarse graining mapping for automatically analyzing protein trajectory dynamics. In particular, we first introduce a graph clustering module that learns to extract coarse grained dynamics by approximating soft spectral clustering. With the clustering assignment function implemented by a graph neural network and parameters learned automatically, we circumvent the need for manual selection of coarse graining mapping. Subsequently, we introduce a path signature transform module served as a feature extractor to characterize the interatomic interactive dynamics after coarse graining. Path signature is a mathematically principled concept that utilizes iterated integrals to describe geometric rough paths in a compact yet rich manner (Lyons, 2014), thus suitable for our tasks where molecular trajectories are highly sampled and non-smooth. After attaching with a task-specific differentiable classifier or regressor, we devise an end-to-end framework, named Deep Signature, for efficiently characterizing the complex protein dynamics. Notably, due to the existence of considerable random fluctuation in simulated trajectories, ideal features ought to maintain symmetry respecting certain geometric transformations. We provide theoretical analysis that our extracted features exhibit invariance to translation, near invariance to rotation, equivariance to permutation of atomic coordinates, and invariance under time reparametrization of paths. Finally, we target our task on predicting functional properties of proteins from MD data, a fundamental task for developing novel drug therapies. We consider three benchmarks including gene regulatory dynamics, epidermal growth factor receptor (EGFR) mutation dynamics, and G protein-coupled receptors (GPCR) dynamics for performance evaluation. The contributions of our paper are as follows:
@@ 2 Related Works
The current investigation into characterizing molecular dynamics, especially interatomic interaction dynamics, remains limited, with only a few studies close to ours. Among them, Endo et al. (2019), Yasuda et al. (2022) and Mustali et al. (2023) are unsupervised methods that build local dynamics ensembles for pre-specified atoms and inspect each atom’s contribution independently. Li et al. (2022) converts MD conformations into images and applies convolutional neural networks to identify diverse active states. Sun et al. (2023) models protein dynamics by representing protein surfaces using implicit neural networks without requiring explicit surface representations. Nevertheless, these approaches cannot capture subtle interatomic interactions along atomic pathways for dissecting protein function, nor is the complexity of their representations independent of time stamps.
@@ 3.2 Deep Signature
$\mathrm{LogSig}^{D}_{r_{i},r_{i+1}}(\tilde{\mathbf{X}})=\sum_{n=1}^{3M}S(\tilde{\mathbf{X}})^{n}_{r_{i},r_{i+1}}e_{n}+\sum_{1\leq n<m\leq 3M}\frac{1}{2}(S(\tilde{\mathbf{X}})^{n,m}_{r_{i},r_{i+1}}-S(\tilde{\mathbf{X}})^{m,n}_{r_{i},r_{i+1}})[e_{n},e_{m}]+\dots$ (10)
@@ 4.1 Experiments on Gene regulatory dynamics
For gene regulatory dynamics, we generate 100 trajectories that describe the interactive dynamics between genes and transcription factors. These dynamics are categorized into degradation type or dimerization type, with an equal number of trajectories for each type. The systems are simulated over a period of $2$s and with a time interval of $0.004$s, resulting in 500 frames per trajectory. Besides, each system encompasses 100 nodes with randomly generated Power-law network structure to describe their relationships. We implement the deep spectral clustering module in our approach using two graph pooling layers that first coarsen the dynamics into 60 nodes and then into 30 nodes. The coefficients of loss terms are set as $\lambda_{1}=1$, $\lambda_{2}=0.01$, and $\lambda_{3}=10$. We optimize our model using Adam with an initial learning rate of 5e-4 and a weight decay of 1e-4. For comparison, we consider several baseline approaches which aggregate nodes’ dynamics by averaging without taking into account their interactions. These methods incorporate a deep spectral clustering model with an MLP to process the first frame only (Head), the last frame only (Tail), the first and last frame (Head & Tail), and all frames with an additional LSTM layer (GraphLSTM) or Transformer (GraphTrans) to aggregate nodes’ temporal dynamics. The overall comparison results with baseline methods are presented in Table 4.
@@
Head 55.28 ±$\pm 10.56$ 30.65 ±$\pm 16.23$
Tail 55.36 ±$\pm 10.72$ 30.71 ±$\pm 16.36$
Head & Tail 67.92 ±$\pm 15.17$ 35.80 ±$\pm 9.27$
GraphLSTM 96.40 ±$\pm 0.89$ 93.14 ±$\pm 1.20$
GraphTrans 86.00 ±$\pm 6.67$ 79.36 ±$\pm 16.37$
Deep Signature 99.12 ±$\pm 0.82$ 98.65 ±$\pm 0.23$
@@ 4.2 Experiments on epidermal growth factor receptor mutation dynamics
$\times$ $\times$ $\checkmark$ 67.60 ±$\pm 6.02$ 19.60 ±$\pm 6.38$
$\times$ $\checkmark$ $\checkmark$ 67.97 ±$\pm 1.15$ 17.60 ±$\pm 4.45$
$\checkmark$ $\times$ $\checkmark$ 66.00 ±$\pm 2.23$ 23.20 ±$\pm 4.99$
$\checkmark$ $\checkmark$ $\checkmark$ 69.33 ±$\pm 4.78$ 21.27 ±$\pm 8.26$
@@ 4.3 Experiments on G protein-coupled receptors dynamics
Accuracy 52.67 ±$\pm 0.94$ 44.00 ±$\pm 1.12$ 53.07 ±$\pm 6.47$ 47.20 ±$\pm 5.86$ 47.20 ±$\pm 5.86$ 53.47 ±$\pm 14.86$ 44.80 ±$\pm 6.97$   58.00 ±$\pm 4.17$
Recall 31.73 ±$\pm 1.77$ 35.47 ±$\pm 3.73$ 26.40 ±$\pm 13.21$ 37.60 ±$\pm 15.72$ 37.60 ±$\pm 15.72$ 38.13 ±$\pm 13.19$ $3$$\pm 21.08$   43.30 ±$\pm 7.85$
The GPCR superfamily is a major therapeutic target as its functioning regulates nearly every physiological process in the human body. To create a dataset for its analysis, we select 26 structures with their MD simulations from the Molecular Dynamics Database for GPCRs (GPCRmd) (Rodríguez-Espigares et al., 2020), including 13 active state structures and 13 inactive state structures. The task is to identify these distinct active states. Each simulation is conducted for 500 ns with a time interval of 200 ps, therefore each trajectory consists of 2,500 frames that describe the conformational dynamics. Here, we only consider atoms in backbones. Experimental results are provided in Table 4.
@@ 5 Conclusion
Acknowledgments
This work was supported in part by the Hong Kong Innovation and Technology Commission (ITC) (InnoHK Project CIMDA), in part by the Institute of Digital Medicine of City University of Hong Kong (Project 9229503), in part by the Hong Kong Research Grants Council under Projects 21200522, 11200323 and 11203220, in part by Chow Sang Sang Donation and Matching Fund (Project 9229161), and in part by the Hong Kong Innovation and Technology Commission (Project GHP/044/21SZ).
@@ B.1 Gene regulatory dynamics
$\frac{\mathrm{d}\bm{x}_{t}(v_{i})}{\mathrm{d}t}=-b_{i}\bm{x}(v_{i})^{f}+\sum_{j=1}^{n}\mathbf{A}^{(i,j)}\frac{\bm{x}^{h}(v_{j})}{\bm{x}^{h}(v_{j})+1},$ (13)
@@ B.3 G protein-coupled receptors dynamics
Appendix C Further Results and Analysis
C.1 Parameter Sensitivity Analysis
The sensitivity of coarsening level. To further validate the robustness of our method for different coarsening levels, we conducted additional experiments by varying the coarsening levels for the EGFR and GPCR datasets. The results are summarized in Table 5.
Table 5: Comparison of EGFR and GPCR Accuracy at Different Coarsening Levels
Coarsening Level 100 50 30 10 5
EGFR Accuracy (%) 66.93
±$\pm 4.62$
69.33
±$\pm 4.78$
66.40
±$\pm 5.06$
63.20
±$\pm 7.56$
57.33
±$\pm 4.73$
GPCR Accuracy (%) 50.67
±$\pm 5.58$
58.00
±$\pm 4.17$
58.00
±$\pm 6.01$
57.73
±$\pm 8.92$
55.07
±$\pm 6.32$
From these results, we observe that coarsening levels of 30 and 50 beads yield the best accuracy for both datasets, demonstrating the robustness of our method across different levels of coarsening. While extreme levels, such as 100 or 5 beads, result in reduced performance, the method is flexible and performs consistently well within a broad range of coarsening levels.
The value of $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$. In the objective of Deep Signature presented in Eq. (12), there exist three hyper-parameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ reweighting the loss terms. We now conduct the parameter sensitivity test by varying their value in $\{10.0,1.0,0.1,0.01\}$. The results are summarized in Table 6, 7 and 8, respectively.
Table 6: Sensitivity analysis for $\lambda_{1}$ and its effect on accuracy.
$\lambda_{1}$ 10.0 1.0 0.1 0.01
Accuracy (%) 59.60
±$\pm 8.34$
69.33
±$\pm 4.78$
67.87
±$\pm 1.95$
67.33
±$\pm 2.77$
Table 7: Sensitivity analysis for $\lambda_{2}$ and its effect on accuracy.
$\lambda_{2}$ 10.0 1.0 0.1 0.01
Accuracy (%) 65.27
±$\pm 2.48$
67.67
±$\pm 1.63$
67.80
±$\pm 1.15$
69.33
±$\pm 4.78$
Table 8: Sensitivity analysis for $\lambda_{3}$ and its effect on accuracy.
$\lambda_{3}$ 10.0 1.0 0.1 0.01
Accuracy (%) 69.33
±$\pm 4.78$
66.53
±$\pm 2.65$
62.00
±$\pm 4.22$
58.93
±$\pm 1.71$
As shown, our model is generally robust to moderate variations in the hyperparameters. For instance, $\lambda_{1}=1.0$, $\lambda_{2}=0.01$, and $\lambda_{3}=10.0$ consistently deliver strong performance. However, extreme values for these parameters (e.g., $\lambda_{3}=0.01$ or $\lambda_{1}=10.0$) can lead to a noticeable drop in accuracy, highlighting the importance of choosing balanced values.